In a test of a gender-selection technique, results consisted of 201 baby girls and 200 baby boys. Based on this result, what is the probability of a girl born to a couple using this technique? Does it appear that the technique is effective in increasing the likelihood that a baby will be a girl? The probability that a girl will be born using this technique is approximately (Type an integer or decimal rounded to three decimal places as needed.)
Answer #1

Number of girls = 201

Number of boys = 200

Total babies = 201+200 = 401

Probability of a girl using this technique = 201/401 = 0.501